OpenStudy (anonymous):

Identify the asymptotes, removable discontinuities, and intercepts for the graph of each function. 1. z(x)= x^2-7x-8 / x^3+64 2. f(x)= 2x^3-3x+1 / x^3 -5x +7

OpenStudy (e.mccormick):

How far did you get with these?

OpenStudy (e.mccormick):

\[z(x)= \frac{x^2-7x-8}{x^3+64}\\ f(x)= \frac{2x^3-3x+1}{x^3 -5x +7}\]

OpenStudy (anonymous):

I dont know how to even start

OpenStudy (e.mccormick):

Factor things. Any factor on the top can be a 0 so an intercept. Any 0 on the bottom can be invalid, so an asymptote. BUT: If a factor is on both the top and bottom it is removable.